### v3

# Duals and time-reversed flows of generalized Ornstein-Uhlenbeck processes

Abstract
We derive explicit representations for the (Siegmund-) dual and the time-reversed flow of generalized Ornstein-Uhlenbeck processes whenever these exist. It turns out that the dual and the process corresponding to the reversed stochastic flow are again generalized Ornstein-Uhlenbeck processes. Further, we observe that the stationary distribution of the dual process provides information about the hitting time of zero of the original process.
AMS 2020 Subject Classifications: 60G51, 60G10, 60H10
Keywords: Generalized Ornstein-Uhlenbeck process, Hitting probability, Ruin probability, Siegmund-duality, Time-reversal
1 Introduction
Almost 50 years ago Siegmund [20] introduced the following duality problem: Given a $[0,\infty]$-valued (universal) time-homogeneous Markov process $(X_{t}^{x})_{t\geq 0}$ (where $X_{0}^{x}=x\geq 0$ denotes the starting value), what are necessary and sufficient conditions for the existence of a dual process, i.e. of another $[0,\infty]$-valued (universal) time-homogeneous Markov process $(Y_{t}^{y})_{t\geq 0}$ (with starting value $Y_{0}^{y}=y\geq 0$) such that $\mathbb{P}(X_{t}^{x}\geq y)=\mathbb{P}(Y_{t}^{y}\leq x)$, i.e.
$\mathbb{P}(X_{t}\geq y|X_{0}=x)=\mathbb{P}(Y_{t}\leq x|Y_{0}=y),\quad 0\leq x,y,t<\infty$ (1.1)
holds?
Note that, similar to the notation in [14, Def. 7.1.1 f.] or [15, Def. 1.21], in the following, we denote by $(X_{t}^{x})_{t\geq 0}$ the time-homogeneous Markov process $(X_{t})_{t\geq 0}$ started in $x$, i.e. $\mathbb{P}(X_{t}^{x}\in B)=\mathbb{P}_{x}(X_{t}\in B)=\mathbb{P}(X_{t}\in B|X_{0}=x)$ for all $B\in\mathcal{B}(\mathbb{R}),t\geq 0$.
Answering above question, Siegmund proved (see [20, Thm. 1]) that for any time-homogeneous Markov process $(X_{t}^{x})_{t\geq 0}$ on $[0,\infty)$ there exists a dual time-homogeneous Markov process $(Y_{t}^{y})_{t\geq 0}$ on $[0,\infty]$ fulfilling (1.1) if and only if there exists some $h>0$ such that
  1. 1.
    the process $(X_{t}^{x})_{t\in[0,h]}$ is stochastically monotone in the sense that the function $x\mapsto\mathbb{P}(X_{t}^{x}\geq y)$ is nondecreasing for any fixed $y\in[0,\infty)$ and $t\in[0,h]$, and
  2. 2.
    the function $x\mapsto\mathbb{P}(X_{t}^{x}\geq y)$ is right-continuous for any fixed $y\in[0,\infty)$ and $t\in[0,h]$.
[…]
Nowadays, the duality relation (1.1) is widely known and commonly used in applied probability. To illustrate its applicability, assume for the moment being that $(Y_{t}^{y})_{t\geq 0}=(R_{t}^{y})_{t\geq 0}$ is a Cramér-Lundberg risk process absorbed in ruin. This process can be defined by
[…]
where
$K_{t}$ $:=ct-\sum_{i=1}^{N_{t}}S_{i},\qquad\tau(y):=\inf\{t\geq 0:y+K_{t}\leq 0\},$
for some premium rate $c\geq 0$, a Poisson process $(N_{t})_{t\geq 0}$ of claim arrivals, and i.i.d. strictly positive claim sizes $\{S_{i},i\in\mathbb{N}\}$ independent of $(N_{t})_{t\geq 0}$. The time $\tau(y)$ is the ruin time. In this case, it is well-known that $(R_{t}^{y})_{t\geq 0}$ is dual in the sense of (1.1) to the M/G/1-queue workload process $(X_{t}^{x})_{t\geq 0}=(V_{t}^{x})_{t\geq 0}$ defined by
$V_{t}^{x}=x+\sum_{i=1}^{N_{t}}S_{i}-\int_{0}^{t}c\mathds{1}_{\{V_{s}^{x}>0\}}\mathop{}\!\mathrm{d}s,\quad x,t\geq 0,$
see [17] for the earliest reference on this relation, or [2, Chapter III.2] for a textbook treatment. (Actually, apart from [1, Example 3.5], where no proof is given, in most references equation (1.1) is only verified for the case $y\geq 0$ and $x=0$. However, the proof given in [2, Chapter III.2] can be easily adapted to work also for general $x,y\geq 0$). In the present situation, the ruin probability can be written as
$\mathbb{P}(\tau(y)<\infty)=\lim_{t\to\infty}\mathbb{P}(R_{t}^{y}\leq 0)=\lim_{t\to\infty}\mathbb{P}(V_{t}^{0}\geq y)=\mathbb{P}(V\geq y),$
where $V$ is a generic random variable whose law is given by the stationary distribution of the dual process $(V_{t})_{t\geq 0}$ which is assumed to exist and to have no atom at $x$, cf. [2, Chapter VII.7].
Thus, in this situation, duality allows to transform the hitting (or ruin) probability of $(R_{t}^{y})_{t\geq 0}$ into the cdf of the stationary distribution of $(V_{t})_{t\geq 0}$, and vice versa. This fact has been used in order to derive eclectic results both in risk theory and queueing theory, see [2, Chapter VII.7] and references therein.
Further examples of pairs of dual Markov processes in the sense of (1.1) can be found e.g. in [1, 10, 21].
Embedding the Cramér-Lundberg process $(R_{t}^{y})_{t\geq 0}$ into an investment market whose dynamics are driven by a Lévy process, say $(W_{t})_{t\geq 0}$, yields a capital process that is often referred to as Paulsen’s risk model [16]. This is a risk process $(R_{t}^{y})_{t\geq 0}$ whose dynamics are described by the stochastic differential equation (SDE)
$\mathop{}\!\mathrm{d}R_{t}^{y}=R_{t-}^{y}\mathop{}\!\mathrm{d}W_{t}+\mathop{}\!\mathrm{d}K_{t},\quad t\geq 0,\quad R_{0}^{y}=y.$ (1.2)
Assuming that $W$ has a jump part of finite variation, in [16] the solution of (1.2) is studied and also an expression for its ruin probability is derived.
In this paper, we treat Siegmund-duality for (time-homogeneous) Markov processes with state space $\mathbb{R}$, defined by (1.1) when $t\geq 0$ and arbitrary $x,y\in\mathbb{R}$ are allowed. We consider solutions of the SDE $\mathop{}\!\mathrm{d}V_{t}^{x}=V_{t-}^{x}\,\mathop{}\!\mathrm{d}U_{t}+\mathop{}\!\mathrm{d}L_{t}$, $V_{0}^{x}=x$, for a general bivariate driving Lévy process $(U_{t},L_{t})_{t\geq 0}$. Assuming that $U$ does not admit jumps of size $-1$ these solutions are known as generalized Ornstein-Uhlenbeck (GOU) processes, see [6, Thm. 2.1], and they can be expressed explicitly. We use this fact to prove in Section 3 that the dual of a GOU process $(V_{t}^{x})_{t\geq 0}$, if existent, is again a GOU process, namely the GOU process $(R_{t}^{y})_{t\geq 0}$ defined by (1.2). We furthermore discuss the relation between hitting probabilities and stationary distributions of the dual pair of GOU processes. As it turns out, the appearing (causal) stationary distribution of the dual GOU process coincides with the non-causal stationary distribution of the original GOU process. When $L$ is a subordinator, then $(V_{t}^{x})_{t\geq 0}$ can also be viewed as a $[0,\infty)$-valued Markov process when only $x\geq 0$ is allowed, while the derived dual GOU process $(R_{t}^{y})_{t\geq 0}$ on all of $\mathbb{R}$ is driven by the negative of a subordinator, hence becomes also negative, even when $y\geq 0$. Hence, the Siegmund-dual process $\widehat{R}_{t}^{y}$ of $(V_{t}^{x})_{t\geq 0}$, $x\geq 0$, viewed as an $[0,\infty]$-valued process, must be different from the Siegmund-dual GOU process $(R_{t}^{y})_{t\geq 0}$, viewed as a Markov process on $\mathbb{R}$. In Proposition 3.9 we clarify the relation between $R$ and $\widehat{R}$.
As it is often the case for dual processes, there is an intrinsic relation between time-reversal of a Markov process (more precisely, the stochastic process induced by the reversed stochastic flow) and its dual process, see e.g. [21, 13], and even the duality relation of the Cramér-Lundberg risk process absorbed in ruin and the M/G/1-queue workload process as described above is proved in [2, Chapter III.2] via a time-reversal argument. Motivated by this, we turn our attention to time-reversed GOU processes in Section 4 and prove that the process associated to the reversed stochastic flow is again a GOU process, and - moreover - that it provides a version of the dual GOU process.
2 Preliminaries
General information regarding Lévy processes can be found e.g. in [7, 8, 19], and for stochastic integration we refer to [18]. Consider the stochastic differential equation (SDE)
$\mathop{}\!\mathrm{d}V_{t}^{x}=V_{t-}^{x}\mathop{}\!\mathrm{d}U_{t}+\mathop{}\!\mathrm{d}L_{t},\;t\geq 0,\quad V_{0}^{x}=x,$ (2.1)
where $(U_{t},L_{t})_{t\geq 0}$ is a bivariate Lévy process with characteristic triplet $(\gamma_{U,L},\Sigma_{U,L},\nu_{U,L})$ for some (standard) location parameter $\gamma_{U,L}\in\mathbb{R}^{2}$, a non-negative definite Gaussian covariance matrix $\Sigma_{U,L}=\left(\begin{smallmatrix}\sigma_{U}^{2}&\sigma_{U,L}\\ \sigma_{U,L}&\sigma_{L}^{2}\end{smallmatrix}\right)\in\mathbb{R}^{2\times 2}$, and a Lévy measure $\nu_{U,L}$ on $\mathbb{R}^{2}\setminus\{0\}$. The characteristic function of $(U_{t},L_{t})$ at $(x,y)^{T}\in\mathbb{R}^{2}$ is then given by
$\mathbb{E}\mathrm{e}^{\mathrm{i}(xU_{t}+yL_{t})}=$ $\exp\Big{\{}t\Big{(}\mathrm{i}(x,y)\gamma_{U,L}-(x,y)\Sigma_{U,L}(x,y)^{T}$
$\quad+\int_{\mathbb{R}^{2}\setminus\{0\}}(\mathrm{e}^{\mathrm{i}(x,y)z}-1-\mathrm{i}(x,y)z\mathds{1}_{|z|\leq 1})\,\nu_{U,L}(\mathop{}\!\mathrm{d}z)\Big{)}\Big{\}}.$
We assume that $\nu_{U,L}(\{-1\}\times\mathbb{R})=0$, i.e. $\Delta U_{t}\neq-1$ for all $t\geq 0$. Hereby, for any càdlàg process $Z$ we denote by $Z_{t-}$ the left-hand limit of $Z$ at time $t\in(0,\infty)$, set $Z_{0-}:=Z_{0}$, and write $\Delta Z_{t}=Z_{t}-Z_{t-}$ for its jumps. The marginal jump measures of $U$ and $L$ are denoted by $\nu_{U}$, and $\nu_{L}$, respectively.
Under this assumption it has been shown in [6, Thm. 2.1] that the solution $(V_{t}^{x})_{t\geq 0}$ of (2.1), the generalized Ornstein-Uhlenbeck (GOU) process, is given by
$V_{t}^{x}=\mathcal{E}(U)_{t}\Big{(}x+\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\mathop{}\!\mathrm{d}\eta_{s}\Big{)}.$ (2.2)
Hereby, $(\mathcal{E}(U)_{t})_{t\geq 0}$ is the stochastic exponential of $(U_{t})_{t\geq 0}$, i.e. the unique solution of the SDE
$\mathop{}\!\mathrm{d}\mathcal{E}(U)_{t}=\mathcal{E}(U)_{t-}\mathop{}\!\mathrm{d}U_{t},\quad\mathcal{E}(U)_{0}=1,$
which can be expressed explicitly by the Doléans-Dade formula (see [18, Thm. II.37]) as
$\mathcal{E}(U)_{t}=\mathrm{e}^{U_{t}-\sigma_{U}^{2}t/2}\prod_{0<s\leq t}(1+\Delta U_{s})\mathrm{e}^{-\Delta U_{s}},\quad t\geq 0,$ (2.3)
from which it follows that $\mathcal{E}(U)_{t}>0$ for all $t$ if and only if $\Delta U_{t}>-1$ for all $t$. The Lévy process $(\eta_{t})_{t\geq 0}$ in (2.2) is given by
$\eta_{t}=L_{t}-\sum_{0<s\leq t}\frac{\Delta U_{s}\Delta L_{s}}{1+\Delta U_{s}}-t\sigma_{U,L},\quad t\geq 0.$ (2.4)
Observe that $\eta_{t}=L_{t}$ if $U$ and $L$ are independent, and that if $L$ is a subordinator and $\Delta U_{t}>-1$ for all $t$, then also $\eta$ is a subordinator, since then $\sigma_{U,L}=0$ and $\Delta\eta_{t}=\Delta L_{t}/(1+\Delta U_{t})$. If $U_{t}=\lambda t$ is chosen deterministically we refer to the resulting GOU process
$V_{t}^{x}=\mathrm{e}^{-\lambda t}\Big{(}x+\int_{(0,t]}\mathrm{e}^{\lambda s}\mathop{}\!\mathrm{d}\eta_{s}\Big{)}$
as Lévy-driven Ornstein-Uhlenbeck process.
Throughout the paper let $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$ denote the augmented natural filtration of $(U_{t},L_{t})_{t\geq 0}$, and note that by [5, Thm. 3.1] (see also [11]), $(V_{t}^{x})_{t\geq 0}$ is a rich Feller process and hence a Markov process with respect to $\mathbb{F}$.
According to [6, Thm. 2.1], in case of its (almost sure) convergence the law of the exponential functional
$\int_{0}^{\infty}\mathcal{E}(U)_{s-}\mathop{}\!\mathrm{d}L_{s}:=\lim_{t\to\infty}\int_{(0,t]}\mathcal{E}(U)_{s-}\mathop{}\!\mathrm{d}L_{s},$
is the unique causal stationary distribution of $(V_{t})_{t\geq 0}$, while the unique, non-causal stationary distribution of $(V_{t})_{t\geq 0}$ is given by the law of
$-\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}:=-\lim_{t\to\infty}\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\,\mathop{}\!\mathrm{d}\eta_{s},$
whenever this converges almost surely. Hereby, a solution of (2.1) with (possibly random) starting point $V_{0}$ is called causal, if $V_{0}$ is independent of $(U,L)$, whereas in case of dependence it is called non-causal. Necessary and sufficient conditions for almost sure convergence of the integrals $\int_{(0,\infty)}\mathcal{E}(U)_{s-}\mathop{}\!\mathrm{d}L_{s}$ and $-\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}$ have been obtained by Erickson and Maller [9] and are summarised in [6, Thm. 3.5 – Cor. 3.7].
3 The dual of the generalized Ornstein-Uhlenbeck process
Under certain conditions the class of Lévy-driven OU processes is self-dual in the following sense:
Consider a storage process solving the SDE
$\mathop{}\!\mathrm{d}V_{t}^{x}=-V_{t}^{x}\mathop{}\!\mathrm{d}t+\mathop{}\!\mathrm{d}L_{t},\;t\geq 0,\quad V_{0}^{x}=x\geq 0,$
where $(L_{t})_{t\geq 0}$ is a compound Poisson process with non-negative jumps. This process can be interpreted as a Lévy-driven Ornstein-Uhlenbeck process with sticky boundary in zero. According to [1, Example 3.5], the dual of the process $(V_{t}^{x})_{t\geq 0}$ is again a Lévy-driven OU process whose driving process is just $(-L_{t})_{t\geq 0}$ and that is killed when passing 00.
In this section we consider arbitrary stochastically monotone GOU processes and show that a similar stability property holds true in a much broader setting.
3.1 Siegmund duality on $\mathbb{R}$
Note that if $\Delta U_{t}>-1$ for all $t$ and $L$ is a subordinator, then $(V_{t}^{x})_{t\geq 0}$ has state-space $[0,\infty)$, and hence this GOU process fits directly into the setting originally considered in [20]. For arbitrary choices of $(U,L)$ this is not the case and we shall therefore first consider an extension of Siegmund’s notion of duality to processes on the whole real line as given in [10, Example 1.1], before we come back to the aforementioned special case of processes on the half-line in Section 3.3.
Definition 3.1.
Let $(X_{t}^{x})_{t\geq 0}$ and $(Y_{t}^{y})_{t\geq 0}$ be (universal) time-homogeneous Markov processes on $\mathbb{R}$, where $x,y\in\mathbb{R}$. Then $(Y_{t}^{y})_{t\geq 0}$ is said to be dual (in the sense of Siegmund) to $(X_{t}^{x})_{t\geq 0}$ if $\mathbb{P}(X_{t}\geq y|X_{0}=x)=\mathbb{P}(Y_{t}\leq x|Y_{0}=y)$ for all $0\leq t<\infty$ and $x,y\in\mathbb{R}$, i.e. if for all $x,y\in\mathbb{R}$
$\mathbb{P}(X_{t}^{x}\geq y)=\mathbb{P}(Y_{t}^{y}\leq x),\quad 0\leq t<\infty.$ (3.1)
Note that it follows immediately from (3.1) that if $(Y_{t}^{y})$ is dual to $(X_{t}^{x})$, then $(-X_{t}^{x})$ is dual to $(-Y_{t}^{y})$. However, duality as defined is not necessarily a symmetric relation as illustrated by the following example.
Example 3.2.
Let $(N_{t})_{t\geq 0}$ be a homogeneous Poisson process (or any other non-trivial subordinator) and define the time-homogeneous Markov processes $X=(X_{t}^{x})_{t\geq 0}$ and $Y=(Y_{t}^{y})_{t\geq 0}$ by
$X_{t}^{x}=\begin{cases}x,&X_{0}^{x}=x\geq 0,\\ x-N_{t},&X_{0}^{x}=x<0,\end{cases}\qquad Y_{t}^{y}=\begin{cases}y,&Y_{0}^{y}=y\geq 0,\\ \min\{y+N_{t},0\},&Y_{0}^{y}=y<0.\end{cases}$
Then it follows immediately that
$\mathbb{P}(X_{t}^{x}\geq y)=\left.\begin{cases}\mathds{1}_{x\geq y},&x\geq 0,\\ \mathbb{P}(N_{t}\leq x-y),&x<0\end{cases}\right\}=\mathbb{P}(Y_{t}^{y}\leq x),$
and hence $(Y_{t}^{y})$ is dual to $(X_{t}^{x})$ in the sense of Definition 3.1. However we note that for $y<0$
$\mathbb{P}(Y_{t}^{y}\geq 0)$ =$\mathbb{P}(y+N_{t}\geq 0)\not=0=\mathbb{P}(X_{t}^{0}\leq y),$
and hence $(X_{t}^{x})$ is not dual to $(Y_{t}^{y})$.
Remark 3.3.
From the definition, one can easily derive conditions for symmetry of the duality relation as follows.
  1. 1.
    Suppose that $(Y_{t}^{y})$ is dual to $(X_{t}^{x})$. Then a necessary and sufficient condition for $(X_{t}^{x})$ to be dual to $(Y_{t}^{y})$ is that
    $\mathbb{P}(X_{t}^{x}=y)\,=\,\mathbb{P}(Y_{t}^{y}=x)\quad\forall\;x,y\in\mathbb{R},\;t\geq 0,$ (3.2)
    as is easily seen from (3.1).
  2. 2.
    Provided that $(Y_{t}^{y})$ is dual to $(X_{t}^{x})$, a sufficient (but not necessary) condition for (3.2) to hold is that $\mathbb{P}(X_{t}^{x}\geq y)$ is continuous as a function in $x$ for each fixed $y\in\mathbb{R}$ and $t>0$, and continuous as a function in $y$ for each fixed $x\in\mathbb{R}$ and $t>0$.
Remark 3.4.
Similar to the original setting of Siegmund-duality described in the introduction, where $[0,\infty]$-valued Markov processes were considered, it is trivial from (3.1) that a necessary condition for an $\mathbb{R}$-valued Markov process $(X_{t}^{x})$ to have a dual process as described in Definition 3.1 is that the function $x\mapsto\mathbb{P}(X_{t}^{x}\geq y)$ is nondecreasing and right-continuous for all $y\in\mathbb{R}$ and $t\geq 0$.
The next theorem shows that the dual of a GOU process $(V_{t}^{x})$ is again a GOU process $(R_{t}^{y})$, and that then $(V_{t}^{x})$ is also dual to $(R_{t}^{y})$. Note that in order to ensure stochastic monotonicity we restrict ourselves to the case $\Delta U_{t}>-1$ for all $t$; see also Remark 3.6 below.
Theorem 3.5.
Let $(V_{t}^{x})_{t\geq 0}$ be a solution of the SDE (2.1) for the bivariate Lévy process $(U_{t},L_{t})_{t\geq 0}$ with $\Delta U_{t}>-1$ for all $t$, as given in (2.2). Then there exists another bivariate Lévy process $(W_{t},K_{t})_{t\geq 0}$ defined by
$\begin{pmatrix}W_{t}\\ K_{t}\end{pmatrix}=\begin{pmatrix}-U_{t}+\sum_{0<s\leq t}\frac{(\Delta U_{s})^{2}}{1+\Delta U_{s}}+t\sigma_{U}^{2}\\ -\eta_{t}\end{pmatrix},\quad t\geq 0,$
such that the unique solution $(R_{t}^{y})_{t\geq 0}$ of
$\mathop{}\!\mathrm{d}R_{t}^{y}=R_{t-}^{y}\mathop{}\!\mathrm{d}W_{t}+\mathop{}\!\mathrm{d}K_{t},\quad R_{0}^{y}=y,$ (3.3)
is dual to $(V_{t}^{x})_{t\geq 0}$ in the sense of Definition 3.1. Moreover, $(V_{t}^{x})_{t\geq 0}$ is dual to $(R_{t}^{y})_{t\geq 0}$.
Proof.
Observe first that $\Delta U_{t}>-1$ for all $t$ implies $\mathcal{E}(U)_{t}>0$ for all $t\geq 0$ by (2.3). Thus, using (2.2),
$V_{t}^{x}\geq y\quad$ $\Leftrightarrow\quad\mathcal{E}(U)_{t}\bigg{(}x+\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\mathop{}\!\mathrm{d}\eta_{s}\bigg{)}\geq y$
$\Leftrightarrow\quad\mathcal{E}(U)_{t}^{-1}y-\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\mathop{}\!\mathrm{d}\eta_{s}\leq x.$ (3.4)
Further, by [6, Lemma 3.4],
$\mathcal{E}(U)_{t}^{-1}=\mathcal{E}(W)_{t},\quad t\geq 0,$ (3.5)
for the Lévy process $(W_{t})_{t\geq 0}$ given by
$W_{t}=-U_{t}+\sum_{0<s\leq t}\frac{(\Delta U_{s})^{2}}{1+\Delta U_{s}}+t\sigma_{U}^{2},$ (3.6)
such that in particular $\Delta W_{t}=\frac{-\Delta U_{t}}{1+\Delta U_{t}}>-1$. Moreover, for every $t\geq 0$, by [6, Lemma 3.1],
[…]
for some Lévy process $(Z_{t})_{t\geq 0}$ that can be computed via [6, Eq. (2.2)] as
$Z_{t}$ $=\eta_{t}-\sum_{0<s\leq t}\frac{\Delta W_{s}\Delta\eta_{s}}{1+\Delta W_{s}}-t\sigma_{W,\eta}$
$=L_{t}-\sum_{0<s\leq t}\frac{\Delta U_{s}\Delta L_{s}}{1+\Delta U_{s}}-t\sigma_{U,L}-\sum_{0<s\leq t}\frac{\frac{-\Delta U_{s}}{1+\Delta U_{s}}\frac{\Delta L_{s}}{1+\Delta U_{s}}}{1+\frac{-\Delta U_{s}}{1+\Delta U_{s}}}+t\sigma_{U,L}=L_{t},$
via (3.6) and (2.4), from which we also derived that $\sigma_{W,\eta}=-\sigma_{U,\eta}=-\sigma_{U,L}$. Thus, inserting (3.5) in (3.4), and applying (3.7) we observe
$\mathbb{P}(V_{t}^{x}\geq y)$ $=\mathbb{P}\bigg{(}\mathcal{E}(W)_{t}y-\int_{(0,t]}\mathcal{E}(W)_{s-}\mathop{}\!\mathrm{d}\eta_{s}\leq x\bigg{)}$
$=\mathbb{P}\bigg{(}\mathcal{E}(W)_{t}y-\mathcal{E}(W)_{t}\int_{(0,t]}\mathcal{E}(W)_{s-}^{-1}\mathop{}\!\mathrm{d}L_{s}\leq x\bigg{)}$
$=\mathbb{P}(R_{t}^{y}\leq x)$
for the GOU process
$R_{t}^{y}=\mathcal{E}(W)_{t}\bigg{(}y+\int_{(0,t]}\mathcal{E}(W)_{s-}^{-1}\mathop{}\!\mathrm{d}(-L_{s})\bigg{)},\quad t\geq 0.$ (3.8)
Defining the auxiliary Lévy process
$-\xi_{t}=\log\mathcal{E}(W)_{t}=W_{t}-\frac{1}{2}\sigma_{W}^{2}t+\sum_{0<s\leq t}\bigg{(}-\Delta W_{s}+\log(1+\Delta W_{s})\bigg{)},\quad t\geq 0,$
using (2.3), we obtain that the SDE solved by $(R_{t}^{y})_{t\geq 0}$ is hence of the form (3.3) with $(W_{t})_{t\geq 0}$ as given in (3.6) and $(K_{t})_{t\geq 0}$ following via [6, Eq. (1.3)] and (3.6) as
$K_{t}$ $=-L_{t}+\sum_{0<s\leq t}(\mathrm{e}^{-\Delta\xi_{s}}-1)(-\Delta L_{s})-t\sigma_{\xi,-L}$
$=-L_{t}+\sum_{0<s\leq t}\frac{\Delta U_{s}\Delta L_{s}}{1+\Delta U_{s}}+t\sigma_{U,L}=-\eta_{t},$
since $\sigma_{\xi,-L}=\sigma_{-W,-L}=\sigma_{U,-L}=-\sigma_{U,L}$. As we have shown that $\mathbb{P}(V_{t}^{x}\geq y)=\mathbb{P}(R_{t}^{y}\leq x)$ for all $x,y\in\mathbb{R}$ and $t\geq 0$, the process $(R_{t}^{y})_{t\geq 0}$ is dual to $(V_{t}^{x})_{t\geq 0}$. Similar calculations, or an application of the obtained result on $(R_{t}^{y})_{t\geq 0}$, show that $\mathbb{P}(V_{t}^{x}=y)=\mathbb{P}(R_{t}^{y}=x)$ so that also $(V_{t}^{x})_{t\geq 0}$ is dual to $(R_{t}^{y})_{t\geq 0}$ by (3.2). ∎
Remark 3.6.
Let $(V_{t}^{x})_{t\geq 0}$ be the GOU process defined by (2.1), where $\Delta U_{t}\neq-1$ for all $t$. Then the condition $\Delta U_{t}>-1$ is not only sufficient for a dual process to $(V_{t}^{x})_{t\geq 0}$ in the sense of Definition 3.1 to exist as seen in the previous theorem, but also necessary. To see this, assume that $U$ has jumps of size less than $-1$ with positive probability. Then from (2.2), for $t>0$,
$\mathbb{P}(V_{t}^{x}\geq y)$ $=\mathbb{P}\bigg{(}\mathcal{E}(U)_{t}\left(x+\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\mathop{}\!\mathrm{d}\eta_{s}\right)\geq y,\mathcal{E}(U)_{t}>0\bigg{)}$
$\quad+\mathbb{P}\bigg{(}\mathcal{E}(U)_{t}\left(x+\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\mathop{}\!\mathrm{d}\eta_{s}\right)\geq y,\mathcal{E}(U)_{t}<0\bigg{)}$
which converges to $\mathbb{P}(\mathcal{E}(U)_{t}>0)$ as $x\to\infty$, and the latter probability is in $(0,1)$ by (2.3). However, if a dual process $(R_{t}^{y})_{t\geq 0}$ as in Definition 3.1 were to exist, then $\lim_{x\to\infty}\mathbb{P}(V_{t}^{x}\geq y)=1$ for all $y\in\mathbb{R}$ by (3.1), since we only allowed $\mathbb{R}$-valued dual processes. Even if we had allowed [,][$\infty$,$\infty$][ - ∞ , ∞ ]-valued dual processes $(R_{t}^{y})_{t\geq 0}$ in Definition 3.1, then still a dual Markov process $(R_{t}^{y})_{t\geq 0}$ would not exist. For if there were such an [,][$\infty$,$\infty$][ - ∞ , ∞ ]-valued Markov process $(R_{t}^{y})$, then $\mathbb{P}(R_{t}^{y}<\infty)=\mathbb{P}(\mathcal{E}(U)_{t}>0)$ for $y\in\mathbb{R}$ by the same argument, hence
$\mathbb{P}(\mathcal{E}(U)_{t}>0)$ $\geq\mathbb{P}(R_{t}^{y}\leq x)=\mathbb{P}(V_{t}^{x}\geq y)=\mathbb{P}\bigg{(}\mathcal{E}(U)_{t}\left(x+\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\,\mathop{}\!\mathrm{d}\eta_{s}\right)\geq y\bigg{)}$
for all $x,y\in\mathbb{R}$, and the latter expression converges to $1$ as $y\to-\infty$, a contradiction.   $\Box$
3.2 On the relation between dual GOU processes and hitting times
As described in the introduction, Siegmund’s duality is often used to match ruin probabilities, i.e. hitting probabilities of the negative half line, with stationary distributions of the dual processes. For the GOU process Paulsen [16, Thm. 3.2] provides an expression for its ruin probability in the case that the driving processes have finite jump activity. His result has been generalized to the setting of this paper in [3, Thm. 2.4] where it reads as follows.
Proposition 3.7.
Consider the GOU process $(V_{t}^{x})_{t\geq 0}$ solving (2.1) for some bivariate Lévy process $(U,L)$ with $\Delta U_{t}>-1$ for all $t$, where $x>0$. Assume that $|\mathcal{E}(U)_{t}|$ converges almost surely to +$\infty$+ ∞ as $t\to\infty$ and that the integral $\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}$ converges almost surely as $t\to\infty$ to a non-degenerate random variable $V_{\infty}=\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}$ with distribution function $H$. Let $\tau(x)=\inf\{t\geq 0,V_{t}^{x}\leq 0\}$. Then
$\mathbb{P}(\tau(x)<\infty)\,\mathbb{E}[H(-V_{\tau(x)}^{x})|\tau(x)<\infty]=H(-x),$
where $\mathbb{E}[\cdot|\tau(x)<\infty]$ is interpreted as 00 if $\mathbb{P}(\tau(x)<\infty)=0$.
Under the conditions and using the same notations as in Theorem 3.5 and Proposition 3.7, we note that
$H(-x)=\mathbb{P}\left(\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}\leq-x\right)=\mathbb{P}\left(\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}K_{s}\geq x\right).$
As mentioned in the Preliminaries, by [6, Thm. 2.1], in case of its convergence the law of the exponential functional
$\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}K_{s}=-V_{\infty},$
is the unique non-causal stationary distribution of $(V_{t}^{x})_{t\geq 0}$, as well as the unique causal stationary distribution of the dual GOU process $(R_{t}^{y})_{t\geq 0}$. Moreover, $R_{t}^{y}$ converges in distribution to $R_{\infty}:=-V_{\infty}$ as $t\to\infty$ for every $y\in\mathbb{R}$. This suggests to interpret duals of GOU processes as a sort of time-reversal and motivates our study of time-reversed GOU processes in the Section 4. Further, these observations lead to the following corollary.
Corollary 3.8.
Under the conditions of Proposition 3.7 the limit of $\mathbb{P}(V_{t}^{x}\leq 0)$ as $t\to\infty$ exists and is equal to $H(-x)=\mathbb{P}(R_{\infty}\geq x)$, i.e.
$\lim_{t\to\infty}\mathbb{P}(V_{t}^{x}\leq 0)=\mathbb{P}(R_{\infty}\geq x)=\mathbb{P}(\tau(x)<\infty)\,\mathbb{E}[H(-V_{\tau(x)}^{x})|\tau(x)<\infty].$
In particular, if $-L$ is a subordinator, then
$\mathbb{P}(\tau(x)<\infty)=\mathbb{P}(R_{\infty}\geq x)=H(-x).$
Proof.
By Theorem 3.5, not only is $(R_{t}^{y})_{t\geq 0}$ dual to $(V_{t}^{x})_{t\geq 0}$, but also $(V_{t}^{x})_{t\geq 0}$ is dual to $(R_{t}^{y})_{t\geq 0}$. Hence $\mathbb{P}(V_{t}^{x}\leq 0)=\mathbb{P}(R_{t}^{0}\geq x)$ and the latter converges by assumption as $t\to\infty$ to $\mathbb{P}(R_{\infty}\geq x)=\mathbb{P}(-V_{\infty}\geq x)=H(-x)$ (recall that $H$ is continuous). The formula for the limit is then immediate from Proposition 3.7.
For the second statement it suffices to assume that $\mathbb{P}(\tau(x)<\infty)>0$. Note that if $-L$ is a subordinator, also $-\eta$ is a subordinator by the observation just below (2.4), which implies that $\int_{0}^{\infty}\mathcal{E}(U)_{s-}^{-1}\,\mathrm{d}\eta_{s}\leq 0$ a.s. and hence $H(0)=1$. Since $V_{\tau(x)}^{x}\leq 0$ on $\{\tau(x)<\infty\}$ by the càdlàg paths of $V^{x}$ this implies $H(-V_{\tau(x)}^{x})=1$ and hence the statement by Proposition 3.7. ∎
3.3 Non-negative Siegmund dual processes
As mentioned before, if $\Delta U_{t}>-1$ for all $t\geq 0$ and $L$ is a subordinator, then $(V_{t}^{x})_{t\geq 0}$ is a time-homogeneous Markov process with state-space $[0,\infty)$, if one restricts to non-negative starting values $x\in[0,\infty)$. Thus, according to Siegmund’s original work, this GOU process has a dual process on $[0,\infty]$ in the sense of (1.1) that will be derived in the next proposition.
Proposition 3.9.
Let $(V_{t}^{x})_{t\geq 0}$ with $x\geq 0$ be a solution of the SDE (2.1) for the bivariate Lévy process $(U_{t},L_{t})_{t\geq 0}$ with $\Delta U_{t}>-1$ for all $t$ and $L$ being a subordinator. Let $(R_{t}^{y})_{t\geq 0}$ with $y\geq 0$ be the solution of (3.3) and set
$\tau_{R}(y):=\inf\{t\geq 0:R_{t}^{y}\leq 0\}.$
Then the process $(\widehat{R}_{t}^{y})_{t\geq 0}$, $y\geq 0$, given by
$\widehat{R}_{t}^{y}:=R_{t}^{y}\mathds{1}_{t<\tau_{R}(y)},$
is a time-homogeneous Markov process on $[0,\infty)$, and it is dual to $(V_{t}^{x})_{t\geq 0}$, $x\geq 0$, in the sense of (1.1). Moreover, we have $\widehat{R}_{t}^{y}=\max\{R_{t}^{y},0\}$ for $t,y\geq 0$.
Proof.
Recall from Theorem 3.5 and its proof that $(R_{t}^{y})_{t\geq 0}$ is dual to $(V_{t}^{x})_{t\geq 0}$ in the sense of (3.1), i.e.
$\mathbb{P}(V_{t}^{x}\geq y)=\mathbb{P}(R_{t}^{y}\leq x),\quad\forall t\geq 0,\;\forall x,y\in\mathbb{R},$
and that $R_{t}^{y}$ is explicitly given by (3.8). As $(R_{t}^{y})_{t\geq 0}$ being a GOU process - is a time-homogeneous Markov process, its killed version $(\widehat{R}_{t}^{y})_{t\geq 0}$ is a time-homogeneous Markov process as well. Moreover, under the given conditions, $\mathcal{E}(W)_{t}=\mathcal{E}(U)_{t}^{-1}>0$ for all $t\geq 0$, and as $L$ is a subordinator it follows from (3.8) that
$y\leq 0\quad\Rightarrow\quad R_{t}^{y}\leq 0,\;\forall t\geq 0.$
Due to time-homogeneity of the GOU process $(R_{t}^{y})_{t\geq 0}$ this implies for any $y\in\mathbb{R}$ that $R_{t}^{y}\leq 0$ for all $t\geq\tau_{R}(y)$, and in particular
$\widehat{R}_{t}^{y}=R_{t}^{y}\mathds{1}_{t<\tau_{R}(y)}=\max\{R_{t}^{y},0\},\quad t\geq 0.$
Thus, for all $t\geq 0$ and all $x,y\geq 0$
$\mathbb{P}(\widehat{R}_{t}^{y}\leq x)$ $=\mathbb{P}(\max\{R_{t}^{y},0\}\leq x)=\mathbb{P}(R_{t}^{y}\leq x)=\mathbb{P}(V_{t}^{x}\geq y),$
which implies the statement. ∎
Remark 3.10.
Observe that although $(V_{t}^{x})_{t\geq 0}$, $x\in\mathbb{R}$, is dual to $(R_{t}^{y})_{t\geq 0}$, $y\in\mathbb{R}$, in the sense of (3.1) by Theorem 3.5, in the situation of Proposition 3.9 the process $(V_{t}^{x})_{t\geq 0}$, $x\geq 0$, is in general not dual in the sense of (1.1) to $(\widehat{R}_{t}^{y})_{t\geq 0}$, $y\geq 0$. This can be seen from
$\mathbb{P}(\widehat{R}_{t}^{0}\geq 0)=1\neq\mathbb{P}\left(\mathcal{E}(U)_{t}\int_{(0,t]}\mathcal{E}(U)_{s-}^{-1}\mathop{}\!\mathrm{d}\eta_{s}\leq 0\right)=\mathbb{P}(V_{t}^{0}\leq 0)\quad\forall\;t>0,$
provided $\eta$ is not the zero-subordinator. Since the function $[0,\infty)\to[0,1]$, $y\mapsto\mathbb{P}(\widehat{R}_{t}^{y}\geq x)$, is nondecreasing and right-continuous for any $t,x\geq 0$, by Siegmund’s result the process $(\widehat{R}_{t}^{y})_{t\geq 0}$, $y\geq 0$, has a dual on $[0,\infty]$, but it is not $(V_{t}^{x})_{t\geq 0}$, $x\geq 0$.
4 The time-reversed flow of the generalized Ornstein-Uhlenbeck process
In contrast to duality which can be considered as a distributional concept, time-reversals of stochastic processes are typically defined pathwise. Therefore we use a different framework in this section than before, and study the GOU process in terms of its stochastic flow and the reversed stochastic flow. Recall the definition of the time-reversal $\widetilde{X}$ of some càdlàg process $X=(X_{s})_{s\in[0,t]}$ as
$\widetilde{X}_{s}=\begin{cases}0,&\text{if }s=0,\\ X_{(t-s)-}-X_{t-},&\text{if }0<s<t,\\ X_{0}-X_{t-},&\text{if }s=t.\end{cases}$ (4.1)
For $t=1$ this coincides with the definition given in [18, Chapter VI.4]. It is well known that when $X$ is a Lévy process, then $(\widetilde{X}_{s})_{s\in[0,t]}$ is also a Lévy process, with the same distribution as $(-X_{s})_{s\in[0,t]}$, see e.g. [7, Lem. II.2], [8, Thm. 11.4] or [19, Prop. 41.8]. Stochastic integrals with respect to time-reversed Lévy processes are then to be seen with respect to their augmented natural filtration, for which the time-reversed Lévy processes are semimartingales.
As before, let $(V_{t}^{x})_{t\geq 0}$ be a solution of the SDE (2.1) for the bivariate Lévy process $(U_{t},L_{t})_{t\geq 0}$. As no stochastic monotonicity is needed in the context of time-reversals, we shall only assume that $\Delta U_{t}\neq-1$ for all $t$.
Theorem 4.1.
Let $(V_{s}^{x})_{s\geq 0}$ be a solution of the SDE (2.1) for the bivariate Lévy process $(U_{s},L_{s})_{s\geq 0}$ with $\Delta U_{s}\neq-1$ for all $s\geq 0$, as given in (2.2) with $\eta$ defined in (2.4), where $x\in\mathbb{R}$. Let $t>0$ be fixed. Denote the time-reversal of $(U,L,\eta)$ on $[0,t]$ by $(\widetilde{U},\widetilde{L},\widetilde{\eta})$, as defined in (4.1) and define
$T_{s}:=\widetilde{U}_{s}+\sigma_{U}^{2}s+\sum_{0<r\leq s}\frac{(\Delta\widetilde{U}_{r})^{2}}{1-\Delta\widetilde{U}_{r}},\quad 0\leq s\leq t.$
Then $(T_{s},\widetilde{L}_{s},\widetilde{\eta}_{s})_{s\in[0,t]}$ is a Lévy process on $[0,t]$ and, almost surely, we have for all $s\in[0,t]$
$V_{(t-s)-}^{x}=\mathcal{E}(T)_{s}\left(V_{t}^{x}+\int_{(0,s]}\mathcal{E}(T)_{u-}^{-1}\,\mathrm{d}\widetilde{L}_{u}\right).$ (4.2)
Further, the stochastic process corresponding to the reversed stochastic flow induced by (4.2), i.e. the process $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, defined by
$R_{s}^{y}:=\mathcal{E}(T)_{s}\left(y+\int_{(0,s]}\mathcal{E}(T)_{u-}^{-1}\mathop{}\!\mathrm{d}\widetilde{L}_{u}\right),$ (4.3)
is the GOU which is the unique solution of the SDE
$\mathop{}\!\mathrm{d}R_{s}^{y}=R_{s-}^{y}\mathop{}\!\mathrm{d}T_{s}+\mathop{}\!\mathrm{d}\widetilde{\eta}_{s},\quad s\in[0,t],\quad R_{0}^{y}=y.$ (4.4)
Proof.
Since $(U,L,\eta)$ is a Lévy process, so is its time-reversal $(\widetilde{U}_{s},\widetilde{L}_{s},\widetilde{\eta}_{s})_{s\in[0,t]}$ on $[0,t]$, hence also $(T_{s},\widetilde{L}_{s},\widetilde{\eta}_{s})_{s\in[0,t]}$ is a Lévy process. Observe that $\Delta\widetilde{U}_{s}=-\Delta U_{t-s}$ for $0<s<t$, and similarly for $L$ and $\eta$. Since $(U,L,\eta)$ almost surely does not jump at the fixed time $t$, we assume that $(\Delta U_{t},\Delta L_{t},\Delta\eta_{t})=0$ and hence $V_{t}^{x}=V_{t-}^{x}$ everywhere. From (2.2) we obtain
$V_{t}^{x}=\frac{\mathcal{E}(U)_{t}}{\mathcal{E}(U)_{t-s}}\left(V_{t-s}^{x}+\int_{(t-s,t]}\left(\frac{\mathcal{E}(U)_{u-}}{\mathcal{E}(U)_{t-s}}\right)^{-1}\mathop{}\!\mathrm{d}\eta_{u}\right).$ (4.5)
Solving this for $V_{t-s}^{x}$ results in
$V_{t-s}^{x}=\frac{\mathcal{E}(U)_{t-s}}{\mathcal{E}(U)_{t}}\left(V_{t}^{x}-\frac{\mathcal{E}(U)_{t}}{\mathcal{E}(U)_{t-s}}\int_{(t-s,t]}\left(\frac{\mathcal{E}(U)_{u-}}{\mathcal{E}(U)_{t-s}}\right)^{-1}\,\mathrm{d}\eta_{u}\right)$
and taking the càdlàg version $V_{(t-s)-}^{x}$ for $s\mapsto V_{t-s}^{x}$ yields
$V_{(t-s)-}^{x}=\frac{\mathcal{E}(U)_{(t-s)-}}{\mathcal{E}(U)_{t}}\left(V_{t}^{x}-\frac{\mathcal{E}(U)_{t}}{\mathcal{E}(U)_{(t-s)}}\int_{[t-s,t]}\left(\frac{\mathcal{E}(U)_{u-}}{\mathcal{E}(U)_{t-s}}\right)^{-1}\,\mathrm{d}\eta_{u}\right).$ (4.6)
From (2.3) we obtain
$\frac{\mathcal{E}(U)_{(t-s)-}}{\mathcal{E}(U)_{t}}$ $=\mathrm{e}^{U_{(t-s)-}-U_{t}+\sigma_{U}^{2}s/2}\prod_{t-s\leq u\leq t}\left((1+\Delta U_{u})\mathrm{e}^{-\Delta U_{u}}\right)^{-1}$
$=\mathrm{e}^{\widetilde{U}_{s}+\sigma_{U}^{2}s/2}\prod_{0<u\leq s}\mathrm{e}^{-\Delta\widetilde{U}_{u}}(1-\Delta\widetilde{U}_{u})^{-1},$
where we used that $\Delta\widetilde{U}_{u}=-\Delta U_{t-u}$ and $\Delta U_{t}=0$. On the other hand, an easy calculation using (2.3) shows that the last expression is equal to $\mathcal{E}(T)_{s}$, so that
$\frac{\mathcal{E}(U)_{(t-s)-}}{\mathcal{E}(U)_{t}}=\mathcal{E}(T)_{s}.$ (4.7)
Since the right-hand sides of both (4.2) and (4.6) define càdlàg processes in $s$, it is enough to show that for given $s$ they agree almost surely (where the exceptional null set may depend on $s$). Since $\eta$ almost surely does not jump at the fixed time $t-s$, the integral over the compact interval $[t-s,s]$ in (4.6) is almost surely equal to the corresponding integral over the half-open interval $(t-s,s]$. Hence, by (4.6) and (4.7), Equation (4.2) will follow if we can show that, almost surely for each fixed $s\in[0,t]$,
$-\frac{\mathcal{E}(U)_{t}}{\mathcal{E}(U)_{t-s}}\int_{(t-s,t]}\left(\frac{\mathcal{E}(U)_{u-}}{\mathcal{E}(U)_{t-s}}\right)^{-1}\,\mathrm{d}\eta_{u}=\int_{(0,s]}\mathcal{E}(T)_{u-}^{-1}\mathop{}\!\mathrm{d}\widetilde{L}_{u}.$ (4.8)
For the treatment of (4.8), define for fixed $s\in[0,t]$
$(U_{u}^{\prime},\eta_{u}^{\prime}):=(U_{u+t-s}-U_{t-s},\eta_{u+t-s}-\eta_{t-s}),\quad u\in[0,s].$
Then $\mathcal{E}(U)_{u+t-s}/\mathcal{E}(U)_{t-s}=\mathcal{E}(U^{\prime})_{u}$ by (2.3) for $u\in[0,s]$ and
$\frac{\mathcal{E}(U)_{t}}{\mathcal{E}(U)_{t-s}}\int_{(t-s,t]}\left(\frac{\mathcal{E}(U)_{u-}}{\mathcal{E}(U)_{t-s}}\right)^{-1}\,\mathop{}\!\mathrm{d}\eta_{u}$ $=\mathcal{E}(U^{\prime})_{s}\int_{(0,s]}\mathcal{E}(U^{\prime})_{u-}^{-1}\mathop{}\!\mathrm{d}\eta_{u}^{\prime}.$
Denote by
$(\widehat{U_{u}^{\prime}},\widehat{\eta_{u}^{\prime}}):=(U_{s-}^{\prime},\eta_{s-}^{\prime})-(U_{(s-u)-}^{\prime},\eta_{(s-u)-}^{\prime}),\quad u\in[0,s],$
the negative of the time reversal of $(U^{\prime},\eta^{\prime})$ at $s$. Then, with [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denoting quadratic covariation,
$\mathcal{E}(U^{\prime})_{s}\int_{(0,s]}\mathcal{E}(U^{\prime})_{u-}^{-1}\mathrm{d}\eta_{u}^{\prime}=\int_{(0,s]}\mathcal{E}(\widehat{U^{\prime}})_{u-}\mathop{}\!\mathrm{d}\widehat{\eta_{u}^{\prime}}+[\mathcal{E}(\widehat{U^{\prime}}),\widehat{\eta}]_{s};$
this is Lemma 6.1 in [12] when all jumps of $U$ are greater than $-1$, and for general $U$ it follows from the proof of Proposition 8.3 in [4] (the left-hand side is the probability limit of the $-B^{\sigma}$ appearing in that proof, while the right-hand side is the probability limit of $A^{\sigma}$). Since $\mathcal{E}(\widehat{U^{\prime}})_{r}=1+\int_{(0,r]}\mathcal{E}(\widehat{U^{\prime}})_{u-}\mathop{}\!\mathrm{d}\widehat{U_{u}^{\prime}}$ we can rewrite the left-hand side of (4.8) as
$-\frac{\mathcal{E}(U)_{t}}{\mathcal{E}(U)_{t-s}}\int_{(t-s,t]}\left(\frac{\mathcal{E}(U)_{u-}}{\mathcal{E}(U)_{t-s}}\right)^{-1}\,\mathrm{d}\eta_{u}$ $=-\int_{(0,s]}\mathcal{E}(\widehat{U^{\prime}})_{u-}\mathop{}\!\mathrm{d}\left(\widehat{\eta_{u}^{\prime}}+[\widehat{U^{\prime}},\widehat{\eta^{\prime}}]_{u}\right)$
$=\int_{(0,s]}\mathcal{E}(-\widetilde{U})_{u-}\mathop{}\!\mathrm{d}\left(\widetilde{\eta}_{u}-[\widetilde{U},\widetilde{\eta}]_{u}\right),$
where in the last line we used that
$(\widehat{U_{u}^{\prime}},\widehat{\eta_{u}^{\prime}})=(U_{s-}^{\prime}-U_{(s-u)-}^{\prime},\eta_{s-}^{\prime}-\eta_{(s-u)-}^{\prime})=(U_{t-}-U_{(t-u)-},\eta_{t-}-\eta_{(t-u)-})=-(\widetilde{U}_{u},\widetilde{\eta}_{u})$
for $u\in[0,s]$. But $\mathcal{E}(-\widetilde{U})_{u}=\mathcal{E}(T)_{u}^{-1}$ as a consequence of (2.3),
$\widetilde{\eta}_{u}=\widetilde{L}_{u}+\sum_{t-u\leq r<t}\frac{\Delta U_{r}\Delta L_{r}}{1+\Delta U_{r}}+\sigma_{U,L}u=\widetilde{L}_{u}+\sum_{0<r\leq u}\frac{\Delta\widetilde{U}_{r}\Delta\widetilde{L}_{r}}{1-\Delta\widetilde{U}_{r}}+\sigma_{U,L}u$ (4.9)
by (2.4) and (4.1), and hence
$\widetilde{\eta}_{u}-[\widetilde{U},\widetilde{\eta}]_{u}=\widetilde{\eta}_{u}-\sigma_{U,\eta}u-\sum_{0<r\leq u}\Delta\widetilde{U}_{r}\Delta\widetilde{\eta}_{r}=\widetilde{L}_{u},$
thus establishing (4.8) and hence (4.2).
For the proof of (4.4), observe that by (2.2), the unique solution to the SDE (4.4) is given by $R_{s}^{y}=\mathcal{E}(T)_{s}(y+\int_{(0,s]}\mathcal{E}(T)_{u-}^{-1}\mathop{}\!\mathrm{d}N_{u})$, where by (2.4) the process $N$ is given by
$N_{s}=\widetilde{\eta}_{s}-\sum_{0<r\leq s}\frac{\Delta T_{r}\Delta\widetilde{\eta}_{r}}{1+\Delta T_{r}}-s\sigma_{T,\widetilde{\eta}},\quad s\in[0,t].$
Inserting (4.9) for $\widetilde{\eta}$ and observing that $\Delta T_{r}=\Delta\widetilde{U}_{r}/(1-\Delta\widetilde{U}_{r})$, $1+\Delta T_{r}=1/(1-\Delta\widetilde{U}_{r})$ and $\Delta\widetilde{\eta}_{r}=\Delta\widetilde{L}_{r}/(1-\Delta\widetilde{U}_{r})$ we see that $N=\widetilde{L}$. This shows that $(R_{s}^{y})_{s\in[0,t]}$ is the unique solution of (4.4). ∎
From  Theorems 3.5 and 4.1 we deduce that if all jumps of $U$ are greater than $-1$, then the process $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, given in (4.4) is Sigmund-dual to $(V_{s}^{x})_{s\in[0,t]}$, $x\in\mathbb{R}$, with the obvious notion of Sigmund-duality for processes defined only on $[0,t]$.
Corollary 4.2.
Let $(V_{s}^{x})_{s\geq 0}$, $x\in\mathbb{R}$, be the solution of the SDE (2.1) for the bivariate Lévy process $(U_{s},L_{s})_{s\geq 0}$ with $\Delta U_{s}>-1$ for all $s\geq 0$. Fix $t>0$ and consider the stochastic process $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, corresponding to the reversed stochastic flow defined in (4.4). Then $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, is Siegmund-dual to $(V_{s}^{x})_{s\in[0,t]}$, $x\in\mathbb{R}$.
Proof.
The processes $(T,\eta)$ appearing in Theorem 4.1 and $(W,K)$ appearing in Theorem 3.5 are given for $s\in[0,t]$ by (compare also (4.9) and (2.4))
$T_{s}$ $=\widetilde{U}_{s}+\sigma_{U}^{2}s+\sum_{0<r\leq s}\frac{(\Delta\widetilde{U}_{r})^{2}}{1-\Delta\widetilde{U}_{r}},$
$\widetilde{\eta}_{s}$ $=\widetilde{L}_{s}+\sigma_{U,L}s+\sum_{0<r\leq u}\frac{\Delta\widetilde{U}_{r}\Delta\widetilde{L}_{r}}{1-\Delta\widetilde{U}_{r}},$
$W_{s}$ $=-U_{s}+\sigma_{U}^{2}s+\sum_{0<r\leq s}\frac{(\Delta U_{r})^{2}}{1+\Delta U_{r}},\quad\mbox{and}$
$K_{s}$ $=-\eta_{s}=-L_{s}+\sigma_{U,L}s+\sum_{0<r\leq s}\frac{\Delta U_{r}\Delta L_{r}}{1+\Delta U_{r}}.$
But since the time-reversed Lévy process $(\tilde{U}_{s},\tilde{L}_{s})_{s\in[0,t]}$ is equal in law to $(-U_{s},-L_{s})_{s\in[0,t]}$ we conclude that $(T_{s},\widetilde{\eta}_{s})_{s\in[0,t]}$ and $(W_{s},K_{s})_{s\in[0,t]}$ are equal in law. The claim then follows from Theorems 3.5 and 4.1. ∎
Remark 4.3.
Corollary 4.2 is not surprising, since reversibility of the stochastic flow is intrinsically related to duality, as discussed in various articles, see e.g. [21]. Also in our case, it is possible to deduce Theorem 3.5 directly from Theorem 4.1. To see this, let $(U,L)$, $V^{x}_{s}$ and $R^{y}_{s}$ be as in Corollary 4.2. For $u\in[0,t]$ denote by $\varphi_{u,t}:\mathbb{R}\to\mathbb{R}$ the stochastic flow which transports $x$ when $V_{u}(\omega)=x$ to $V_{t}(\omega)$, i.e. $\varphi_{u,t}(x)=V_{t}(x)$ conditional on $V_{u}=x$ (we suppress the superscript $x$ here and prefer to work with $V_{t}|V_{0}=x$ rather than $V_{t}^{x}$ for the moment). By time-homogeneity of $V$, we have for $s\in[0,t]$
$\mathbb{P}(V_{s}\geq y|V_{0}=x)=\mathbb{P}(V_{t}\geq y|V_{t-s}=x)=\mathbb{P}(\varphi_{t-s,t}(x)\geq y)=\mathbb{P}(\varphi_{t-s,t}^{-1}(y)\leq x),$
since $\varphi_{t-s,t}$ is strictly increasing and bijective (the exact form of the flow can be read off from (4.5)). But $\varphi_{t-s,t}^{-1}(y)=R_{s-}^{y}$ by (4.2) and (4.3), hence
$\mathbb{P}(V_{s}\geq y|V_{0}=x)=\mathbb{P}(R_{s-}^{y}\leq x)=\mathbb{P}(R_{s}^{y}\leq x)=\mathbb{P}(R_{s}\leq x|R_{0}=y),$
showing that $(R_{s}^{y})_{s\in[0,t]}$, $y\in\mathbb{R}$, is dual to $(V_{s}^{x})_{s\in[0,t]}$, $x\in\mathbb{R}$. By the proof of Corollary 4.2, $(T_{s},\widetilde{\eta}_{s})_{s\in[0,t]}$ is equal in law to $(W_{s},K_{s})_{s\in[0,t]}$, showing that the process given in Theorem 3.5 (call it $\overline{R}_{s}^{y}$ for the moment) is also dual to $R$. Hence we have given another proof of the duality of $(\overline{R}_{s}^{y})$ to $(V_{s}^{x})$ in Theorem 3.5. This last proof is in line with the reasoning given in Sigman and Ryan [21, Cor. 3.1 (1)], who consider $[0,\infty)$-valued time-homogeneous Markov processes with certain properties.